# How simple can you go? An off-the-shelf transformer approach to molecular dynamics

### v2

@@ Abstract
Most current neural networks for molecular dynamics (MD) include physical inductive biases, resulting in specialized and complex architectures. This is in contrast to most other machine learning domains, where specialist approaches are increasingly replaced by general-purpose architectures trained on vast datasets. In line with this trend, several recent studies have questioned the necessity of architectural features commonly found in MD models, such as built-in rotational equivariance or energy conservation. In this work, we contribute to the ongoing discussion by evaluating the performance of an MD model with as few specialized architectural features as possible. We present a recipe for MD using an Edge Transformer, an “off-the-shelf” transformer architecture that has been minimally modified for the MD domain, termed MD-ET. Our model implements neither built-in equivariance nor energy conservation. We use a simple supervised pre-training scheme on $\sim$30 million molecular structures from the QCML database. Using this “off-the-shelf” approach, we show state-of-the-art results on several benchmarks after fine-tuning for a small number of steps. Additionally, we examine the effects of being only approximately equivariant and energy conserving for MD simulations, proposing a novel method for distinguishing the errors resulting from non-equivariance from other sources of inaccuracies like numerical rounding errors. While our model exhibits runaway energy increases on larger structures, we show approximately energy-conserving NVE simulations for a range of small structures.
@@ 1 Introduction
Molecular dynamics (MD) simulations are essential for understanding molecular behavior [1, 2, 3], such as molecular relaxation [4], predicting structure [5], modeling interactions [6], replicating spectra [7], up to protein folding and design [8, 9, 10]. Machine Learning (ML) models can be used as heuristic approximations of the Schrödinger equation thereby significantly accelerating accurate MD simulations (e.g. Rupp et al. [11], Noé et al. [12], Unke et al. [13], Keith et al. [14], Unke et al. [15]).To ensure physical consistency of predictions, researchers have focused on incorporating inductive biases into MD models, which restrict models to physically plausible solutions [16, 17, 18, 19, 20, 13]. For instance, rotational equivariance guarantees that if a molecule is rotated, the forces are rotated as well (e.g. Schütt et al. [21], Bronstein et al. [22], Unke et al. [23]). Meanwhile, many other ML domains are increasingly using unconstrained, general-purpose architectures and large amounts of training data [24, 25, 26].
@@
Several recent theoretical [27, 28] and empirical works [29, 30, 31, 32] discuss the necessity of commonly used physical constraints and inductive biases in MD models, most notably SO(3)-equivariance and energy conservation. For example, architectures predicting non-conservative forces include ForceNet [33], GemNet-dT [34], Orb [35] or Equiformer [36]. However, to the best of our knowledge, no architectures that fully embrace the general-purpose paradigm have been presented so far. We contribute by studying an extreme case: We present and evaluate the best MD model we can build using a minimal amount of MD-specific design features. Specifically, we aim to use an architecture that could be used in some other ML setting with minimal changes.
@@
The “off-the-shelf” recipe we use is simple: We modify an Edge Transformer (ET) [37, 38] with MD-specific embedding layers and train on the new QCML database [39]. QCML has an unprecedented size and range across the periodic table, including out-of-equilibrium structures, different spin and charge states, and includes properties for a subset of $\sim$30 million entries calculated with density functional theory accuracy at the PBE0 level [40, 41] (including dispersion corrections [42, 43]). During training, we use two randomly rotated and mirrored copies of each sampled structure, so that our model can learn approximate O(3)-equivariance. Instead of predicting a combined loss [44] of energies and forces or calculating forces as the (negative) gradient of the predicted energy with respect to positions, we predict forces directly. This approach does not guarantee that the predicted forces are energy-conserving, but it improves the speed of the model, because forces need not be calculated by automatic differentiation. For downstream task evaluation, we fine-tune the model for a small number of steps only.
For MD simulations, we evaluate either the pre-trained model without further training (zero-shot) or fine-tune on a small number of samples of the target structure (few-shot). To increase the accuracy during MD simulations, we use a frame-averaging approach [45] over the SO(3) group to achieve an approximate “post-hoc” equivariance and further correct the predicted forces to guarantee no spurious rigid rotational or translational motion is introduced over time.
Our contributions are as follows:
@@
    We present an MD-adapted Edge Transformer (MD-ET) and show that our implementation can achieve competitive performance on several common benchmarks. We also demonstrate sample-efficiently finetuning.
@@
    We examine the effects of forgoing almost all commonly used inductive biases. Specifically, we evaluate the effects of being only approximately equivariant and energy-conserving in detail. We present a novel evaluation framework that is able to quantify deviations from true equivariance and distinguish them from effects stemming from model inaccuracy and numerical noise.
@@
    We present a comparison of MD-ET with an identically pretrained equivariant and energy-conserving model, highlighting the impact of removing physical constraints on MD simulations.
@@ 2 Background
While incorporating physics-inspired inductive biases into MLFF models yields powerful and data-efficient architectures, they also impose constraints. For instance, directly predicting (non-conservative) rotationally equivariant forces restricts a model to only performing equivariant operations [46]. A more subtle downside of specialized architectures is the increased difficulty to transfer advances from other ML research and profit from improved hardware and software tools. Engineering complexity was recently cited as the reason the AlphaFold project removed its equivariant inductive biases [10, 8]. Since protein conformer prediction is thematically close to MD, this prompted the discussion of which inductive biases are essential to the MD domain: Questioning the need for equivariance, Langer et al. [47] find that data augmentation during training can achieve a high degree of approximate equivariance in molecular dynamics. Neumann et al. [35] present Orb, a completely unconstrained message-passing architecture with outstanding benchmark performance. Elhag et al. [31] show that an additional loss term to induce equivariance combined with data augmentation results in competitive performance and approximate equivariance. Several theoretical works also suggest that unconstrained models can produce equivariant outputs under certain circumstances Gerken and Kessel [27], Nordenfors and Flinth [28], Puny et al. [45]. In contrast, the performance effects of removing energy conservation are less clear. Although benchmarking results suggest non-conservative models can perform stable MD simulations [35]. Bigi et al. [29] report that all non-conservative models in their study suffer from runaway temperature increases and thus instability.
@@
Several current debates in the MD field center around competing paradigms – one leveraging inductive biases for sample efficiency, and another prioritizing simplicity, scale and transfer using general-purpose models. Recently, several aspects of the scaling paradigm have been studied and discussed in isolation [47, 29, 32]. Our work advances this discourse by empirically evaluating a simple baseline approach that fully adopts the scaling paradigm for MD and exploring its limits.
@@ 3 Molecular Dynamics Edge Transformer
We use the transformer, an architecture which has proven to scale well in many domains, as a starting point in our design process. We decide to use the Edge Transformer variant introduced by Bergen et al. [37] for algorithmic reasoning and adapted by Müller et al. [38] for graph learning, employing a higher-order attention mechanism instead of self-attention due to its expressivity (see below). We modify only the ET’s embeddings for MD. We therefore name our variant MD-ET.
@@
Expressivity describes the ability of an ML model to tell the difference between two non-equivalent inputs, i.e., to distinguish two graphs that are not isomorphic. Morris et al. [48] show that the expressivity of a graph learning model can be related to performing a variant of the Weisfeiler-Leman (WL) graph isomorphism test. Hordan et al. [49] extend the WL-test to equivariant point clouds and show that a cutoff-free model with 3-WL expressivity is universal on 3D point clouds, i.e., such a model can theoretically learn to distinguish any two 3D point clouds. Müller et al. [38] show the ET has an expressive power equivalent to 3-WL. The ET can thus universally distinguish 3D point clouds, making it theoretically well-suited as an MLFF.
@@
The ET derives its name from its tokenization scheme. We represent a molecular system as the set of all edges of a fully connected graph, that is, all pairs of atoms (including self-loops). We then encode the molecular system as a three-dimensional tensor $\boldsymbol{X}\in\mathbb{R}^{N\times N\times D}$, where $N$ denotes the number of atoms in the system and $D$ is the embedding dimension. An embedding $S_{ij}$ thus represents the edge between atoms $i$ and $j$, while $S_{ii}$ represents the atom $i$.
@@
To achieve the expressive power of 3-WL, the ET performs updates on the three-dimensional tensor representation described in the previous paragraph using a triangular attention mechanism, performing interactions between triples of atoms,
$\operatorname{TRIA}(\boldsymbol{x}_{ij})=\sum_{l=1}^{N}\alpha_{ilj}\boldsymbol{V}_{ilj}\ .$ (1)
Triangular attention is defined as a tensor product between an attention tensor $\boldsymbol{\alpha}\in\mathbb{R}^{N\times N\times N}$ and a value tensor $\boldsymbol{V}\in\mathbb{R}^{N\times N\times D}$. An element of the attention tensor is calculated as
@@
is a combination of the value vectors of $\boldsymbol{x}_{il}$ and $\boldsymbol{x}_{lj}$ through elementwise multiplication. $\boldsymbol{W}^{Q}$,$\boldsymbol{W}^{K}$,$\boldsymbol{W}^{V_{1}}$, $\boldsymbol{W}^{V_{2}}$ $\in\mathbb{R}^{d\times d}$ are learned weight matrices [38]. The triangular attention mechanism extends to multi-head attention analogous to regular self-attention. Similarly, a full ET layer is defined analogously to a transformer layer [50]; specifically,
@@
The triangular attention is straightforward to implement and JIT-compile (see Appendix A.2). The large 3D tensor products are embarrassingly parallelizable and thus execute very efficiently on accelerators. This makes the ET comparatively fast despite a runtime and memory complexity of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in the system (see Appendix A.3).
@@
To generate initial edge representations, we combine several embeddings: i) spin and charge, ii) atomic numbers , iii) pairwise distances, and iv) pairwise displacement vectors. To maintain MD-ET’s simplicity we implement commonly used embedding layer where possible. For details, see Appendix A.2.1.
@@ 4 Experiments
To contribute to the current discussion about inductive biases in MD, we need to thoroughly evaluate MD-ET. Specifically, we want to address the following questions:
  • Q1: Can an “off-the-shelf” model without inductive biases compete on common benchmarks measuring accuracy, inference speed, and MD stability?
  • Q2: How equivariant is MD-ET? Is approximate equivariance likely to impact MD simulation stability?
  • Q3: Is MD-ET approximately energy-conserving? How affected is the stability of MD simulations?
  • Q4: How faithful are MD simulations? Is there a correlation between energy conservation and observables?
4.1 Pretraining
We use a pretrained model for all downstream evaluations. However, instead of the commonly used denoising task [52, 35] we perform supervised pretraining on the new QCML dataset [39]. We create an approximate 90%/5%/5% split from QCML. As QCML contains multiple conformations for each structure sampled along their normal modes, we ensure that all conformations of a structure get assigned to the same split.
To learn approximate equivariance, we use data augmentation: During pretraining we duplicate each batch once and apply random rotations and reflections to both copies to form an augmented batch. We train MD-ET for 880k steps on QCML using a batch size of 1024 (i.e., 512 before data augmentation), which takes approximately 16 A100 GPU-days (see Appendix A.5).
4.2 Postprocessing
Our model does not require any modifications for downstream fine-tuning tasks. For MD-simulations we apply two techniques to enhance model predictions:
Net Force and Torque Removal
We adopt the net force and torque removal algorithm described in [35]. We find that removing net forces and torque from model predictions improves the stability of MD simulations, while only increasing the time of a model evaluation by 10-20 percent (see Appendix A.3). We use net force and torque removal in all simulation experiments.
Frame-averaging
For some simulations, we observe that frame-averaging [45] helps to improve MD stability. For frame-averaging we use $M$ rotation matrices to pass $M$ different orientations of the input positions to the model. We then apply the inverse rotations to the predicted forces and average them. We discuss the effects of frame-averaging in section 4.4. We use frame-averaging only where explicitly stated. For implementation details, see section 4.4 and Appendix A.5.
4.3 Q1: Benchmark Results
We evaluate the performance of our proposed MD-ET model across several standard molecular dynamics benchmarks. We compare MD-ET with established MLFFs, focusing on their accuracy, stability and efficiency. Our experiments cover a range of datasets, including QCML [39], MD17 [16], and Ko2020 [53], each presenting distinct challenges.
Performance on the QCML Dataset
Table 1: Force prediction test MAE on the QCML dataset in kcal/mol Å-1 for different models.

### v3

@@ Abstract
Most current neural networks for molecular dynamics (MD) include physical inductive biases, resulting in specialized and complex architectures. This is in contrast to most other machine learning domains, where specialist approaches are increasingly replaced by general-purpose architectures trained on vast datasets. In line with this trend, several recent studies have questioned the necessity of architectural features commonly found in MD models, such as built-in rotational equivariance or energy conservation. In this work, we contribute to the ongoing discussion by evaluating the performance of an MD model with as few specialized architectural features as possible. We present a recipe for MD using an edge transformer (ET), an “off-the-shelf” transformer architecture that has been minimally modified for the MD domain, termed MD-ET. Our model implements neither built-in equivariance nor energy conservation. We use a simple supervised pretraining scheme on $\sim$30 million molecular structures from the QCML database. Using this “off-the-shelf” approach, we show state-of-the-art results on several benchmarks after fine-tuning for a small number of steps. Using MD-ET as a simple and expressive testbed, we examine the effects of being only approximately equivariant and energy conserving for MD simulations and thereby try to evaluate the practical usefulness of unconstrained MD models. While our model exhibits runaway energy increases on larger structures, we show approximately energy-conserving NVE simulations for a range of small structures.
@@ 1 Introduction
Molecular dynamics (MD) simulations are essential for understanding molecular behavior [23, 70, 47], such as molecular relaxation [64], predicting structure [29], modeling interactions [13], replicating spectra [16], up to protein folding and design [35, 75, 1]. Machine Learning (ML) models can be used as heuristic approximations of the Schrödinger equation thereby significantly accelerating accurate MD simulations (e.g. [63, 55, 38, 73]). To ensure physical consistency of predictions, researchers have focused on incorporating inductive biases into MD models, which restrict models to physically plausible solutions [15, 66, 65, 69, 14, 74]. For instance, rotational equivariance guarantees that if a molecule is rotated, the forces are rotated as well (e.g. [67, 10, 5]). In contrast to this, many other ML domains are increasingly using unconstrained, general-purpose architectures and large amounts of training data [36, 21, 17].
@@
Several recent theoretical [30, 56] and empirical works [7, 8, 76, 20, 9] discuss the necessity of commonly used physical constraints and inductive biases in MD models, most notably SO(3)-equivariance and energy conservation. For example, architectures predicting non-conservative forces include ForceNet [34], GemNet-dT [27], Orb [54, 62] or Equiformer [43].
Physical inductive biases dramatically simplify the problem of learning molecular force fields by reducing the dimensionality of the functional solution space and are thus easy to motivate. There are also, however, several motivations to remove them: Common ways to implement equivariance such as group convolutions or irreducible representations are typically computationally more expensive and drastically restrict the design of the neural architecture [18]. Furthermore, there is some theoretical evidence that equivariant networks have inferior training dynamics [56]. Meanwhile deriving energy-conserving forces using automatic differentiation increases wall time by 2–3x, more when considering that automatic optimization frameworks such as torch.compile do not natively support second derivatives as required to train conservative force fields [82], though framework-specific workarounds exist [68]. Furthermore, the need to compute second derivatives only allows the use of components that are smooth everywhere, which, for instance, some activation functions are not. Finally, there are several ML subfields such as graph learning and ML on point clouds which are concerned with problems closely related to MD. If physical inductive biases could be relaxed or approximated sufficiently closely by a general-purpose model these fields could be unified [59].
We thus study an extreme case: We present and evaluate the best MD model we can build using a minimal amount of MD-specific design features. Specifically, we aim to use an architecture that could be used in some related ML setting with minimal changes. Our aim is to explore potential benefits of a fully unconstrained approach to MD, while highlighting problems arising from a lack of physical inductive biases.
@@
Our “off-the-shelf” recipe is simple: We modify an edge transformer (ET) [6, 51] with MD-specific embedding layers and train on the new QCML database [26]. QCML ranges across the periodic table, including out-of-equilibrium structures, different spin and charge states, and includes properties for a subset of $\sim$30 million entries calculated with density functional theory accuracy at the PBE0 level [58, 2] (including dispersion corrections [12, 32]). During training, we use two randomly rotated and mirrored copies of each sampled structure, so that our model can learn approximate O(3)-equivariance. Instead of predicting a combined loss [46] of energies and forces or calculating forces as the negative gradient of the predicted energy with respect to positions, we predict forces directly. For downstream task evaluation, we fine-tune the model for a small number of steps only.
For MD simulations, we evaluate either the pretrained model without further training (zero-shot) or fine-tune using a small number of samples of the target structure (few-shot). Our contributions are as follows:
@@
    We present an MD-adapted edge transformer (MD-ET) and show that our implementation can achieve competitive performance on several common benchmarks. We also demonstrate sample-efficient fine-tuning.
@@
    We examine the effects of forgoing almost all commonly used inductive biases. Specifically, we evaluate the effects of being only approximately equivariant and energy-conserving in detail.
@@
    We present a comparison of MD-ET with an identically pretrained equivariant and energy-conserving model as well as the recently released Orb-v2 model, highlighting the impact of removing physical constraints on MD simulations.
@@ 2 Background
While incorporating physics-inspired inductive biases into machine learning force fields (MLFF) models yields powerful and data-efficient architectures, they also impose constraints. For instance, directly predicting (non-conservative) rotationally equivariant forces restricts a model to only performing equivariant operations [18]. A more subtle downside of specialized architectures is the increased difficulty to transfer advances from other ML research and profit from improved hardware and software tools. Engineering complexity was recently cited as the reason the AlphaFold project removed its equivariant inductive biases [1, 35]. Since protein conformer prediction is thematically close to MD, this prompted discussions on which inductive biases are essential to the MD domain: Questioning the need for equivariance, Langer et al. [42] find that data augmentation during training can achieve a high degree of approximate equivariance in MD. Neumann et al. [54] present Orb, a completely unconstrained message-passing architecture with outstanding benchmark performance and it has also been used in combination with quantum-mechanical analysis for orbital localization in polymer nanocomposites [22]. Elhag et al. [20] show that an additional loss term to induce equivariance combined with data augmentation results in competitive performance and approximate equivariance. Wang et al. [78] demonstrate the potential benefits of unconstrained approaches in molecular conformer generation, achieving state-of-the-art results without relying on domain-specific assumptions. To demonstrate that methods designed to process point clouds can be used for MD, Pozdnyakov and Ceriotti [59] introduce PET, a message-passing architecture where each message-passing block uses a transformer to process local neighborhoods. Specifically, for each atom-centered environment, tokens encoding neighbor information (positions, species, incoming messages) undergo self-attention, with output tokens serving as outgoing messages. While the locality of neighborhoods gives PET linear runtime, self-attention over all neighbor tokens is costly since the same atom pairs contribute to multiple overlapping neighborhoods. PET-based potentials have already been used in materials simulations, e.g., to study reconstructed the $\beta$-lithium thiophosphate surface, $\beta$-Li3PS4 [71]. Furthermore, PET-MAD [48] was developed as a generally applicable interatomic potential trained on a dataset designed to increase atomic diversity [49]. Qu and Krishnapriyan [61] propose a scalable attention-based MLFF, EScAIP, which improves computational efficiency and model expressivity by leveraging attention mechanisms and optimized GPU kernels, achieving state-of-the-art performance across various chemical datasets. The same group has proposed a graph-less model, exploring even further reduction of inductive bias [41]. Several theoretical works also suggest that unconstrained models can produce equivariant outputs under certain circumstances [30, 56, 60]. In contrast, the performance effects of removing energy conservation are less clear. Although benchmarking results suggest non-conservative models can perform stable MD simulations [54], Bigi et al. [7] report that all non-conservative models in their study suffer from runaway temperature increases and thus instability.
@@
Current debates in the field of ML for MD center around competing paradigms—one leveraging inductive biases for sample efficiency, and another prioritizing simplicity, scale and transfer using general-purpose models. Recently, several aspects of the scaling paradigm have been studied and discussed in isolation [42, 7, 9, 79, 24]. Our work advances this discourse by empirically evaluating a simple baseline approach that fully adopts the scaling paradigm for MD and exploring its limits.
@@ 3 Molecular dynamics edge transformer
We use the transformer, an architecture which has proven to scale well in many domains, as a starting point in our design process. Since we want to keep the architecture as close as possible to a regular transformer implementation we also choose to forgo graph construction and thus cutoffs completely. Note that any attention mechanism can be converted into an equivalent sparse message passing algorithm by only evaluating the dot products selectively. Any attention mechanism converted in such a manner will also exhibit a linear runtime. We thus do not see a higher than linear-scaling runtime as an issue for this analysis, even though it is prohibitive for many typical use cases of MLFFs.
We choose to represent an atomic system using its own frame of reference, i.e., using pairwise distances. While this represents a significant inductive bias, it removes several symmetries from the representation without constraining the model architecture. To enable the model to learn equivariant functions we also supply pairwise directional embeddings using absolute polar and azimuthal angles.
Expressivity, as employed in the context of graph learning, is the ability of an ML model to differentiate two non-equivalent inputs, i.e., to distinguish two graphs that are not isomorphic. Morris et al. [50] show that the expressivity of a graph learning model can be related to performing a variant of the Weisfeiler–Leman (WL) graph isomorphism test. Additionally more expressive models can learn more complex features in each layer, which implies that even if a less expressive model can still theoretically learn to distinguish two inputs it might have more difficulty converging to a representation which separates them given the same number of layers [51].
Hordan et al. [33] show that using pairwise distance features alone a 3-WL neural network can distinguish any two euclidean point clouds. Since we use pairwise distance embeddings we thus believe that 3-WL expressivity is a feature for which it is worthwhile to deviate from the simplest possible transformer implementation. The Edge Transformer (ET) variant introduced by Bergen et al. [6] uses a triangular self-attention implementation which is similar to standard dot-product self-attention (see Appendix A.3) and has been shown to be 3-WL expressive [51]. Since it uses a mixture of dot-product attention and sum-pooling it also features a runtime and memory complexity of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms, whereas full edge-to-edge self-attention has a complexity of $\mathcal{O}(E^{2})=\mathcal{O}(N^{4})$ without cutoffs. Merely using atom-to-atom self-attention could further reduce complexity to $\mathcal{O}(N^{2})$, atom-wise tokenization however only allows the use of edge features as (learned) biases in the attention operation, a restriction which we find too limiting. We thus choose to use ET [6] for this study. As noted above we use the "off-the-shelf" dense implementation of ET, which is sufficiently fast for molecules smaller than approximately 100 atoms (see Appendix A.4) and merely add embedding layers for MD. We therefore name our variant MD-ET.
@@
The ET derives its name from its tokenization scheme. We represent a molecular system as the set of all edges of a fully connected graph, that is, all pairs of atoms (including self-loops). We then encode the molecular system as a three-dimensional tensor $\boldsymbol{x}\in\mathbb{R}^{N\times N\times D}$, where $N$ denotes the number of atoms in the system and $D$ is the embedding dimension. An embedding $\boldsymbol{x}_{ij}$ thus represents the edge between atoms $i$ and $j$, while $\boldsymbol{x}_{ii}$ represents the atom $i$.
@@
To achieve the expressive power of 3-WL, the ET performs updates on $\boldsymbol{x}$ using a triangular attention mechanism, performing interactions between triples of atoms,
$\operatorname{TRIA}(\boldsymbol{x}_{ij})=\sum_{l=1}^{N}\alpha_{ilj}\boldsymbol{v}_{ilj}\ ,$ (1)
a tensor product between an attention tensor $\boldsymbol{\alpha}\in\mathbb{R}^{N\times N\times N}$ and a value tensor $\boldsymbol{v}\in\mathbb{R}^{N\times N\times D}$. An element of the attention tensor is calculated as
@@
is a combination of the value vectors of $\boldsymbol{x}_{il}$ and $\boldsymbol{x}_{lj}$ through elementwise multiplication $\odot$. $\boldsymbol{W}^{Q}$$\boldsymbol{W}^{K}$$\boldsymbol{W}^{V_{1}}$, $\boldsymbol{W}^{V_{2}}$ $\in\mathbb{R}^{D\times D}$ are learned weight matrices [51]. The triangular attention mechanism extends to multi-head attention analogous to regular self-attention. Similarly, a full ET layer is defined analogously to a transformer layer [77]; specifically,
@@
To generate initial edge representations, we combine several embeddings: i) spin and charge, ii) atomic numbers, iii) pairwise distances, and iv) pairwise displacement vectors. To maintain MD-ET’s simplicity, we implement commonly used embedding layers where possible. For details, see Appendix A.2.
Limitations
@@ 4 Experiments
To contribute to the current discussion of inductive biases in MD, we need to thoroughly evaluate MD-ET. Specifically, we want to address the following questions:
  • Can an “off-the-shelf” model without inductive biases compete on common benchmarks measuring accuracy, inference speed, and MD stability?
  • How equivariant is MD-ET? Are there systematic biases depending on system orientation?
  • Is MD-ET approximately energy-conserving? How stable are NVE simulations?
  • How faithful are NVT simulations? Are observables impacted by non-conservative/non-equivariant force predictions?
We use a pretrained model for all downstream evaluations. However, instead of the commonly used denoising task [80, 54], we perform supervised pretraining on the new QCML dataset [26]. We create an approximate 90%/5%/5% split from QCML. As QCML contains multiple conformations for each structure sampled along their normal modes, we ensure that all conformations of a structure get assigned to the same split. To learn approximate equivariance, we use data augmentation: During pretraining we duplicate each batch once and apply random rotations and reflections to both copies to form an augmented batch. We train MD-ET for 880k steps on QCML using a batch size of 1024 (i.e., 512 before data augmentation), which takes approximately 16 A100 GPU-days (see Appendix A.5). Our model does not require any modifications for downstream fine-tuning tasks. For MD-simulations we adopt the net force and torque removal algorithm described in [54]. We find that removing net forces and torque from model predictions improves the stability of MD simulations while only increasing the time of a model evaluation by 10-20 percent (see Appendix A.4). We use net force and torque removal in all simulation experiments.
4.1 Unconstrained models are competitive on common MD benchmarks
We evaluate the performance of our proposed MD-ET model across several standard MD benchmarks. We compare MD-ET with established MLFFs, focusing on their accuracy, stability, and efficiency. We evaluated MD-ET on QCML [26], MD17 [15], Ko2020 [39], xxMD [57] and SPICE [19], each presenting distinct challenges.
Pretraining performance
Table 1: Force prediction test MAE on the QCML dataset in meV Å-1 for different models.